# Health Text Simplification: An Annotated Corpus for Digestive Cancer Education and Novel Strategies for Reinforcement Learning

### v2

@@ 3.1.2 Med-EASi
Our experimentation utilized the Med-EASi corpus, which is a crowd-sourced medical text simplification corpus [37]. It includes annotations for specific simplification edits, including changes associated with elaboration, replacement, deletion, and insertion. The introduction of Med-EASi includes simplification results based on fine-tuning T5-large [46]. Experimentation included a conventional machine translation approach using T5, as well as a controlled simplification approach that utilized these edit labels (elaboration, replacement, deletion, and insertion). Med-EASi uses a novel annotation approach that combines the efforts of medical experts, layperson crowd-workers, and AI algorithms. The dataset includes 1,979 expert-layperson sentence pairs, covering a wide range of medical topics and complexities. The average FKGL of the expert and layperson samples is approximately 13th and 10th grades, respectively.
@@ 3.2 Simplification Models
Our exploration of SimpleDC and Med-EASi focused on state-of-the-art encoder-only LLMs, including Llama 2 [13] and GPT-4 [14]. Using Llama 2, we explored zero-shot and SFT approaches to establish performance baselines and introduce a novel RL approach for simplification. Using GPT-4, we explored different prompt-based strategies and presented a novel self-correction strategy for simplification.
@@ 3.2.1 Supervised fine-tuning (SFT)
As a baseline, we trained Llama 2 on original-simplified sentence pairs using conventional SFT with cross-entropy loss. SFT experimentation included SimpleDC and Med-EASi.
Figure 2: Reinforcement Learning with Human Feedback for Medical Text Simplification
@@ 3.2.2 Reinforcement Learning
We explored RL strategies to incorporate unlabeled in-domain text into model training. Using Proximal Policy Optimization (PPO) [59], we investigated RL reward functions that include metrics related to reading level assessment (ie, FKGL), topical relevance (semantic similarity), and a data-driven approach to identify simplified text (referred to as Original vs. Simplified below). This exploration includes Reinforcement Learning with Human Feedback (RLHF) [60]. Figure 2 presents an overview of the RL pipeline. Each reward metric is described below. RL frameworks commonly use reward values in the range [0, 1], where higher values indicate more desirable outcomes and lower values indicate less desirable outcomes. For each reward metric presented, individual and composite rewards are scaled to this range.
Readability Reward: FKGL is a heuristic function to assess reading level. Although an imperfect metric, FKGL provides an automatic way to assess the reading grade level of the generated text. The TESLEA approach [38] includes a normalization strategy to convert FKGL scores to the range [-1,1], based on the target reading level. To map the FKGL scores to the range [0, 1], we took inspiration from the TESLEA approach. The reading level reward, $R_{FKGL}$, was calculated by normalizing the FKGL score by the reading level target and applying a sigmoid function:
@@
Relevance Reward: The relevance reward, $R_{Rel}$, is intended to quantify the semantic relationship between the simplified and original texts, where a higher reward indicates greater semantic similarity. We follow the TESLEA approach [38] for relevance scoring: 1) the original and simplified texts are mapped to separate vector representations using an encoder and 2) semantic similarity is assessed based on cosine similarity between these two vectors. BioSentVec [61] was selected as the encoder, as it has been extensively trained on biomedical literature. Cosine similarity falls in the range [0, 1], where higher values indicate higher semantic correspondence. The relevance reward incentivizes the preservation of semantic relatedness between the original and simplified texts. We used the pre-trained BioSentVec without any training on SimpleDC.
Original vs. Simplified (OvS) Reward: We introduce a new RL reward, $R_{OvS}$, based on the likelihood that a sample has been simplified. Using SimpleDC training data, we trained a binary classifier to label samples as original or simplified using BioMed-RoBERTa base [62]. This reward function captures some aspects of the annotators’ (human) preferences and is considered RLHF. The OvS classification model achieved an accuracy of 75.1% on the SimpleDC development data set. This performance demonstrates the ability to identify key features of the original and simplified texts. The reward value, $R_{OvS}$, is the predicted probability that the text is simplified ($P(simplified)$) providing a continuous reward signal ranging from [0, 1], where higher values suggest better simplification. Our intuition behind this reward function is that it provides a mechanism for learning features of the original and simplified text and propagating this learning to unlabeled samples during RL.
@@
    $\mathbf{R_{FKGL+Rel}}$: We combined the reading level reward, $R_{FKGL}$, and the relevance reward, $R_{Rel}$, to create an aggregated reward, $R_{FKGL+Rel}=H(R_{FKGL},R_{Rel})$. This aggregated reward is intended to capture both text complexity and semantic similarity to yield text that is at the desired reading level, while preserving the original meaning. This reward function does not incorporate any learning from SimpleDC and is therefore considered RL.
@@
    $\mathbf{R_{FKGL+OvS}}$: We also combined the reading level reward, $R_{FKGL}$, and the Original vs. Simplified reward, $R_{OvS}$, to create an aggregated reward, $R_{FKGL+OvS}=H(R_{FKGL},R_{OvS})$. This aggregated reward is intended to assess text complexity ($R_{FKGL}$) and identify linguistic features that differentiate the original and simplified texts ($R_{OvS}$) to produce text that is both at the desired reading level and emulates the simplification of the SimpleDC. As this reward function incorporates a training signal derived from the SimpleDC annotations, it is considered RLHF.
In our initial trials, we explored weighted averages, drawing inspiration from the TESLEA paper [38]. However, due to inconsistent outcomes with hyperparameters and lower performance in general, we have chosen to focus on experiments employing the harmonic mean in our presentation.
@@ 3.2.3 Fine-tuning With Reward Models
    RL/RLHF only: A pre-trained LLM was trained using RL/RLHF using unlabeled in-domain data (digestive cancer text) to learn the target domain and the text simplification task without any SFT of the LLM on annotated in-domain data (SimpleDC). The goal of this exploration was to understand the achievable performance without any SFT.
@@
In this experimentation, the unlabeled in-domain text consisted of 1,000 web pages from ACS, CDC, and NCI related to digestive cancer that were not annotated and that did not overlap with SimpleDC.
In all RL experiments, the training used PPO, where a Kullback-Leibler (KL) divergence constraint [63] was implemented to limit excessive deviation of the RL-trained LLM from the original LLM. This divergence constraint provides a regulatory mechanism to ensure the LLM maintains linguistic coherence and avoids overfitting to potentially incorrect or excessive scores from the reward model. This constraint was essential due to the inherent limitations of the reward model, which may not accurately capture the quality of the generated text.
@@ 3.2.4 Prompt-based models
Given our focus on developing simplification models with limited training data and the success of GPT-4 in many tasks, we present prompt-based approaches for text simplification with GPT-4 [14], including in-context learning, CoT, and self-correction.
@@
FKGL-Enhanced Prompt: Building on the initial prompt, we added a detailed description of the FKGL heuristic, indicating that lower reading levels are associated with shorter words and sentences, and provided the FKGL formula for calculating the grade level from word and sentence lengths.
@@
Chain-of-Thought (CoT): Our CoT approach builds on the FKGL-Enhanced Prompt approach by providing the model with additional sample-specific context, including the average word length, average sentence length, and a list of longer words ($\geq$ 3 syllables) for each sample. This model aims to provide explicit guidance for the simplification process and anchor the simplification in commonly used readability standards. Figure 3 illustrates the CoT approach.
@@
Self-Correction (SC): Inspired by recent work on self-correction [51, 52], we created the self-correction approach for simple text simplification shown in Figure 4. The intent was to both exploit the reasoning capabilities of the LLM and incorporate additional external learning signals to iteratively refine the output. Each generated simplification was scored for reading level using FKGL and semantic fidelity using BERTScore. Acceptance criteria include: 1) FKGL score $\leq 6$ or FKGL reduced by 3 grade levels from the original text and 2) BERTScore between the generated and original texts $\geq 0.95$. If the simplification met the acceptance criteria, it was accepted; otherwise, the simplification was fed back into GPT-4 with instructions for improving the simplification. Each sample was processed iteratively at most 3 times.
@@ 3.2.5 Experimental Paradigm
Our experimental pipelines utilized PyTorch [64] and Hugging Face’s Transformers [65] package. We used two Llama 2 variants: Llama-2-7b-chat-hf  [^2] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf and Llama-2-7B-32K-Instruct  [^3] https://huggingface.co/togethercomputer/Llama-2-7B-32K-Instruct. Training and evaluation utilized A100 GPUs. We also used Weights and Biases [66] for tracking training and performance. For SFT, we utilized LORA [67] utilities from the Parameter-Efficient Fine-Tuning (PEFT) package [68, 69], to reduce the time and space complexity of model training. Hyperparameters were selected based on computational constraints and performance optimization. We selected a batch size of 8 to balance memory usage and model updating. The learning rate was set to 2e-4, following preliminary experiments to find a balance between catastrophic forgetting and suboptimal convergence. Additional hyperparameters, such as weight decay (0.001) and maximum gradient norm (0.3), were included to prevent overfitting and maintain the stability of the training process. For RLHF, we specified a range of hyperparameters, including a learning rate of 1.41e-5 and PPO epochs of 2, reflecting a conservative approach to incremental learning. Our choice of batch and mini-batch sizes, set to 4, was influenced by the need to manage the computational load effectively while still capturing the nuances of gradient updates during training.
@@ 3.3 Evaluation
We utilized widely accepted evaluation metrics, each addressing specific aspects of text simplification quality. These include: 1) System output Against References and against the Input sentence (SARI) [70], which evaluates the quality of simplification by considering additions, deletions, and substitutions; 2) Bilingual Evaluation Understudy (BLEU) [71, 72], which evaluates the precision of the generated text against reference translations; 3) BERTScore [73], which quantifies the semantic similarity between texts using contextual embeddings; 4) Recall-Oriented Understudy for Gisting Evaluation (ROUGE) [74], focusing n-gram overlap between the output and reference texts; and 5) Flesch-Kincaid Grade Level (FKGL) [58], measuring the readability of the text based on sentence and word length.
@@ 4 Results
SARI BLEU
@@
Score
ROUGE FKGL
@@
Llama-2-7B- 32K-Instruct Zero-shot 03 27.21 0.07 0.87 0.31 9.13
SFT 04 47.82 0.47 0.95 0.70 7.04
Llama-2-7b- chat-hf RL ($R_{FKGL+Rel}$) 05 29.99 0.29 0.91 0.54 8.29
RLHF ($R_{OvS}$) 06 31.62 0.30 0.91 0.56 8.40
RLHF ($R_{FKGL+OvS}$) 07 29.03 0.28 0.91 0.54 8.38
SFT+RLHF ($R_{FKGL+OvS}$) 08 60.39 0.70 0.97 0.89 8.24
Llama-2-7B- 32K-Instruct RL ($R_{FKGL+Rel}$) 09 43.47 0.54 0.94 0.72 8.24
RLHF ($R_{OvS}$) 10 45.09 0.40 0.92 0.61 8.05
RLHF ($R_{FKGL+OvS}$) 11 59.02 0.48 0.93 0.65 7.54
SFT+RLHF ($R_{FKGL+OvS}$) 12 74.18 0.77 0.97 0.88 7.95
GPT4 Turbo Zero-shot 13 22.03 0.14 0.92 0.43 5.71
In-context 14 22.04 0.14 0.92 0.42 5.68
FKGL-Enhanced 15 23.78 0.24 0.94 0.53 5.37
CoT 16 20.97 0.18 0.93 0.46 4.71
Self-Correction 17 22.86 0.21 0.93 0.51 4.84
@@
Model Name RLHF Reward
@@
SARI BLEU
@@
Score
ROUGE FKGL
Llama-2-7b- chat-hf none 18 35.15 0.52 0.94 0.64 8.49
$R_{FKGL}+R_{Rel}$ 19 45.71 0.63 0.96 0.78 7.50
$R_{FKGL}+R_{OvS}$ 20 74.41 0.73 0.97 0.89 7.74
Llama-2-7B- 32K-Instruct none 21 26.70 0.36 0.92 0.52 8.25
$R_{FKGL}+R_{Rel}$ 22 37.55 0.58 0.95 0.72 7.45
$R_{FKGL}+R_{OvS}$ 23 45.71 0.63 0.96 0.78 7.50
Table 2: In-domain and domain adaptation performance of Prompt-based, SFT, and RLHF experimentation on SimpleDC dataset. Source FKGL: 8.11, Target FKGL: 7.65.

### v3

@@ 3.1.2 Med-EASi
Our experimentation utilized the Med-EASi [43] corpus, which includes simplifications for medical texts annotated through a combination of expert, layperson, and AI-generated contributions. The more complex medical texts are annotated by experts, while simpler texts are handled by layperson crowd-workers, often assisted by AI-generated suggestions. It includes annotations for specific edit types: elaboration, replacement, deletion, and insertion. There are 1,979 original-simplified sentence pairs, covering a wide range of medical topics. The average FKGL of the original and simplified samples is approximately 13th and 10th grades, respectively. The introduction of Med-EASi includes simplification results based on fine-tuning T5-large [52]. Experimentation included a conventional machine translation approach using T5, as well as a controlled simplification approach that utilized the edit types.
@@ 3.2 Simplification Models
Our exploration of SimpleDC and Med-EASi focused on state-of-the-art decoder-only LLMs, including Llama 2 [14], Llama 3 [15] and GPT-4 [16]. Using Llama models, we explored zero-shot, SFT, and RL/RLHF approaches. Using GPT-4, we explored prompt-based strategies, including a novel self-correction strategy.
@@ 3.2.1 Supervised fine-tuning (SFT)
As a baseline, we trained Llama models on original-simplified sentence pairs using conventional SFT with cross-entropy loss. SFT experimentation included SimpleDC and Med-EASi.
Figure 2: Reinforcement Learning from Human Feedback for Medical Text Simplification
@@ 3.2.2 Reinforcement Learning
We explored RL strategies to incorporate unlabeled in-domain text into training. Using Proximal Policy Optimization (PPO) [68], we investigated RL reward functions that include metrics related to reading level assessment (i.e., FKGL), topical relevance (semantic similarity), and a data-driven approach to identify simplified text (referred to as Original vs. Simplified below). This exploration includes RLHF [69]. Figure 2 presents an overview of the RL pipeline, each reward metric is described below. RL frameworks commonly use reward values in the range $[0,1]$, where higher values indicate more desirable outcomes and lower values indicate less desirable outcomes. For each reward metric presented, individual and composite rewards are scaled to this range.
Readability Reward: FKGL is a heuristic function to assess reading level. Although an imperfect metric, FKGL provides an automatic way to assess the reading grade level of the text. The TESLEA approach [44] includes a normalization strategy to convert FKGL scores to the range $[-1,1]$, based on the target reading level. Our reading level reward, $R_{FKGL}$, which was inspired by the TESLEA approach, was calculated by normalizing the FKGL score by the reading level target and applying a sigmoid function:
@@
Relevance Reward: The relevance reward, $R_{Rel}$, is intended to quantify the semantic relationship between the simplified and original texts, where a higher reward indicates greater semantic similarity. We follow the TESLEA approach [44] for relevance scoring: 1) the original and simplified texts are mapped to separate vector representations using an encoder and 2) semantic similarity is assessed based on cosine similarity between these vectors. BioSentVec [70] was selected as the encoder, as it has been extensively trained on biomedical literature. Cosine similarity falls in the range $[0,1]$, where higher values indicate higher semantic correspondence. The relevance reward incentivizes the preservation of semantic relatedness between the original and simplified texts. We used the pre-trained BioSentVec without any training on SimpleDC.
Original vs. Simplified (OvS) Reward: We introduce a new RL reward, $R_{OvS}$, based on the likelihood that a sample has been simplified. Using SimpleDC training data, we trained a binary classifier to label samples as original or simplified using BioMed-RoBERTa base [71]. Similar to InstructGPT [72], we utilize a classifier trained on human-annotated data; however, in our formulation, the classifier distinguishes between original and simplified texts. This classifier provides a probabilistic reward signal reflecting human judgment, capturing annotators’ preferences and is considered RLHF. The OvS classifier achieved 75.1% accuracy on the SimpleDC development set. This performance demonstrates the ability to identify key features of the original and simplified texts. The reward value, $R_{OvS}$, is the predicted probability that the text is simplified ($P(simplified)$) providing a continuous reward signal ranging from $[0,1]$, where higher values suggest better simplification. Our intuition behind this reward function is that it provides a mechanism for learning features of the original and simplified text and propagating this learning to unlabeled samples during RL.
@@
    $\mathbf{R_{FKGL+Rel}}$: We combined the reading level reward, $R_{FKGL}$, and the relevance reward, $R_{Rel}$, to create an aggregated reward, $R_{FKGL+Rel}=H(R_{FKGL},R_{Rel})$. This reward is intended to capture both text complexity and semantic similarity to yield text that is at the desired reading level and preserves the original meaning. It does not incorporate any learning from SimpleDC and is therefore considered RL.
@@
    $\mathbf{R_{FKGL+OvS}}$: We also combined the reading level reward, $R_{FKGL}$, and the Original vs. Simplified reward, $R_{OvS}$, to create an aggregated reward, $R_{FKGL+OvS}=H(R_{FKGL},R_{OvS})$. This aggregated reward is intended to assess text complexity ($R_{FKGL}$) and identify linguistic features that differentiate the original and simplified texts ($R_{OvS}$) to produce text that is both at the desired reading level and emulates the simplification of SimpleDC. As this reward function incorporates a training signal derived from the SimpleDC annotations, it is considered RLHF.
In initial experimentation, we explored weighted averages, drawing inspiration from the TESLEA paper [44]. However, due to inconsistent outcomes with hyperparameters and lower performance, we chose to focus on experiments employing the harmonic mean.
@@ 3.2.3 Fine-tuning With Reward Models
    RL/RLHF only: A pre-trained LLM was trained using RL/RLHF with unlabeled in-domain data (digestive cancer text) to learn the target domain and simplification task without any SFT of the LLM on annotated in-domain data (SimpleDC). The goal of this exploration was to understand the achievable performance without any SFT.
@@
In this experimentation, the unlabeled in-domain text consisted of 1,000 web pages from ACS, CDC, and NCI related to digestive cancer that were not annotated and did not overlap with SimpleDC.
In all RL experiments, training used PPO, where a Kullback-Leibler (KL) divergence constraint [73] was implemented to limit excessive deviation of the RL-trained LLM from the original LLM. This divergence constraint provides a regulatory mechanism to ensure the LLM maintains linguistic coherence and avoids overfitting to potentially incorrect or excessive scores from the reward model. This constraint was essential due to the inherent limitations of the reward model, which may not accurately capture the quality of the generated text.
@@ 3.2.4 Prompt-based models
Given our focus on developing simplification models with limited training data and the success of GPT-4 in many tasks, we explored prompt-based approaches for text simplification with GPT-4 [16], including in-context learning, CoT, and self-correction.
@@
FKGL-Enhanced Prompt: Building on the initial prompt, we added a detailed description of the FKGL heuristic, indicating that lower reading levels are associated with shorter words and sentences, and provided the FKGL formula for grade-level calculations.
@@
Chain-of-Thought (CoT): Our CoT approach builds on the FKGL-Enhanced Prompt approach by providing the model with additional sample-specific context, including the average word length, average sentence length, and a list of longer words ($\geq 3$ syllables) for each sample. This model aims to provide explicit guidance for the simplification process and anchor the simplification in commonly used readability standards. Figure 3 illustrates the CoT approach.
@@
Self-Correction (SC): Inspired by recent work on self-correction [57, 58], we created the self-correction approach for simplification shown in Figure 4. The intent was to both exploit the reasoning capabilities of the LLM and incorporate additional external learning signals to iteratively refine the output. Each generated simplification was scored for reading level using FKGL and semantic fidelity using BERTScore. Acceptance criteria include: 1) FKGL score $\leq 6$ or FKGL reduced by three grade levels from the original text and 2) BERTScore between the generated and original texts $\geq 0.95$. If the simplification met the acceptance criteria, it was accepted; otherwise, the simplification was fed back into GPT-4 with instructions for improving the simplification. Each sample was processed iteratively at most three times.
@@ 3.2.5 Experimental Paradigm
We utilized PyTorch [74] and Hugging Face’s Transformers [75]. We used one Llama 2 variant: Llama-2-7b-chat-hf  [^3] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf and two Llama 3 variants: Llama-3-8B-Instruct  [^4] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct and Llama-3-70B-Instruct  [^5] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct. Experimentations were performed using NVIDIA A100 GPUs.
For the Llama-2-7b-chat-hf model, all experiments, including zero-shot, SFT, and RL/RLHF, were conducted using a single A100-80GB GPU. For the Llama-3-8B-Instruct model, zero-shot and SFT experiments were also feasible using a single A100 GPU; however, RL/RLHF experimentation required two A100-80GB GPUs. For the Llama-3-70B-Instruct model, zero-shot inference was conducted using four A100-80GB GPUs without quantization; however, SFT required 4-bit quantization [76] with four A100-80GB GPUs. Unfortunately, RL/RLHF experimentation with Llama-3-70B-Instruct was not feasible with our computational resources, even with the 4-bit quantization and utilization of four A100-80GB GPUs.
We also used Weights and Biases [77] for tracking training and performance. For SFT, we utilized LORA [78] utilities from the Parameter-Efficient Fine-Tuning (PEFT) package [79, 80], to reduce the time and space complexity of model training. Hyperparameters were selected based on computational constraints and performance optimization. We selected a batch size of 8 to balance memory usage and model updating. The learning rate was set to 2e-4, following preliminary experiments to find a balance between catastrophic forgetting and suboptimal convergence. Additional hyperparameters, such as weight decay (0.001) and maximum gradient norm (0.3), were included to prevent overfitting and maintain the stability of the training process. For RLHF, we specified a range of hyperparameters, including a learning rate of 1.41e-5 and PPO epochs of 2, reflecting a conservative approach to incremental learning. Our choice of batch and mini-batch sizes, set to 4, was influenced by the need to manage the computational load effectively while still capturing the nuances of gradient updates during training.
@@ 3.3 Evaluation
We utilized widely accepted evaluation metrics, each addressing specific aspects of text simplification quality, including:
  1. 1.
    System output Against References and against the Input sentence (SARI) [81]: SARI evaluates simplification quality by comparing system-generated output with both the original and reference texts. Outputs are assessed based the operations: add - words that are in the reference but not in the original text; keep - words from the original text are also in the reference, and delete - words from the original text that are not in the reference.
  2. 2.
    Bilingual Evaluation Understudy (BLEU) [82, 83]: BLEU measures the precision of n-grams in the generated text against a reference by evaluating how many words in the generated text are also present in the reference text. We used 4-gram BLEU.
  3. 3.
    BERTScore [84]: BERTScore quantifies the semantic similarity between texts using contextual embeddings from the BERT model. It compares sentence-level embeddings of the generated text and reference texts, providing a measure of semantic comparison beyond surface-level text matching.
  4. 4.
    Recall-Oriented Understudy for Gisting Evaluation (ROUGE) [85]: ROUGE focuses on n-gram overlap between the output and reference texts. ROUGE-N measures the overlap of n-grams, and ROUGE-L measures the longest common subsequence. We used ROUGE-L because it is less sensitive to minor differences in wording and is particularly useful for assessing fluency and coherency.
  5. 5.
    Flesch-Kincaid Grade Level (FKGL) [67]: FKGL assesses the reading level of text based on the average number of syllables per word and number of words per sentence. FKGL scores indicate the estimated grade level.
@@ 4 Results
SARI
($\uparrow$)
BLEU
($\uparrow$)
@@
Score ($\uparrow$)
ROUGE
($\uparrow$)
FKGL
($\downarrow$)
@@
Llama-3- 8B-Instruct Zero-shot 03 25.15 0.11 0.90 0.46 8.96
SFT 04 52.04 0.42 0.93 0.61 6.30
Llama-3- 70B-Instruct Zero-shot 05 24.49 0.02 0.84 0.12 8.59
SFT 06 46.73 0.39 0.93 0.56 5.16
Llama-2-7b- chat-hf RL ($R_{FKGL+Rel}$) 07 29.99 0.29 0.91 0.54 8.29
RLHF ($R_{OvS}$) 08 31.62 0.30 0.91 0.56 8.40
RLHF ($R_{FKGL+OvS}$) 09 29.03 0.28 0.91 0.54 8.38
SFT+RLHF ($R_{FKGL+OvS}$) 10 60.39 0.70 0.97 0.89 8.24
Llama-3- 8B-Instruct RL ($R_{FKGL+Rel}$) 11 20.44 0.11 0.88 0.32 7.38
RLHF ($R_{OvS}$) 12 42.45 0.37 0.91 0.51 8.02
RLHF ($R_{FKGL+OvS}$) 13 55.01 0.59 0.93 0.80 8.24
SFT+RLHF ($R_{FKGL+OvS}$) 14 50.98 0.53 0.91 0.73 8.11
GPT-4 Turbo Zero-shot 15 22.03 0.14 0.92 0.43 5.71
In-context 16 22.04 0.14 0.92 0.42 5.68
FKGL-Enhanced 17 23.78 0.24 0.94 0.53 5.37
CoT 18 20.97 0.18 0.93 0.46 4.71
Self-Correction 19 22.86 0.21 0.93 0.51 4.84
@@
Model Name Variant
@@
SARI
($\uparrow$)
BLEU
($\uparrow$)
@@
Score ($\uparrow$)
ROUGE
($\uparrow$)
FKGL
($\downarrow$)
Llama-2-7b- chat-hf none 20 35.15 0.52 0.94 0.64 8.49
$R_{FKGL}+R_{Rel}$ 21 45.71 0.63 0.96 0.78 7.50
$R_{FKGL}+R_{OvS}$ 22 74.41 0.73 0.97 0.89 7.74
Llama-3- 8B-Instruct none 23 38.73 0.43 0.93 0.60 6.36
$R_{FKGL}+R_{Rel}$ 24 49.97 0.45 0.92 0.59 7.27
$R_{FKGL}+R_{OvS}$ 25 51.06 0.57 0.91 0.72 8.39
Table 3: In-domain and domain adaptation performance of Prompt-based, SFT, and RLHF experimentation on SimpleDC dataset. Source FKGL: 8.11, Target FKGL: 7.65. Higher scores indicate better performance for SARI, BLEU, BERTScore, and ROUGE ($\uparrow$), whereas lower scores are preferred for FKGL ($\downarrow$).